# itemizestditemize

### v2

@@ I Introduction
Figure 1: Scenario of Collaborative Road Detection with 45 Aerial Robots. The drones are deployed in an unknown environment and tasked to collaboratively detect the roads and visually cover them. Particularly, the drones aim to collectively maximize the total covered road area in their top-view field of view (FOV). At each motion-planning step, each drone chooses a few other drones to coordinate with, subject to its onboard communication and computation bandwidth constraints. The full collaborative autonomy pipeline is depicted in Fig. 2. (a)–(c) depict the task progress across time: (a) The drones start from their initial deployment points; (b) Progress after 3 replanning steps; (c) The drones have achieved maximal road coverage.
@@
The current approaches are either near-optimal but not scalable, or real-time but offer no performance guarantees. The near-optimal methods [9, 10, 11] that are being used for active information gathering with multiple robots [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20], scale cubically with the network size, resulting in impractical running times in mesh networks with tens of robots (up to tens of minutes per action coordination step in the presence of real-world communication delays). The state-of-the-art multi-robot motion planning methods are real-time [21, 22, 23, 24] but they are suitable in spatial exploration where robots are expected to operate far from each other. Then, the submodular structure can be ignored and opportunistic coordination is sufficient. We discuss the related work in extension in Section II.
@@
  • Time complexity: RAG’s decision-time scales linearly with the network size. Instead, the competitive near-optimal submodular optimization algorithms scale cubically with the network size. To this end, RAG introduces a resource-minimal distributed optimization paradigm that requires each robot to make decisions based only on locally observed information and information received from and about their neighbors only. In contrast, the current submodular optimization paradigms [9, 10, 11], widely applied in robotics and control [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20], that instead require the relay of information about all robots across the network.
@@ II Related Work
The current approximation paradigms for submodular maximization problem [9, 10, 11] have been widely used in active information gathering with multiple robots [12, 2, 13, 14, 1, 15, 16, 17, 18, 19, 20] (see also the survey [5]). However, they may not scale. The reason is that their communication and computation complexities are superlinear in the number of robots, particularly, quadratic or cubic (Table I). Instead, we provide an algorithm with linear communication and computation complexity.
@@
The quadratic or higher complexity of [9, 10, 11] over r2r networks is due to their coordination protocols: they instruct the robots to retain and relay information about all or most other robots in the network. Particularly, [9, 10] require iterative decision-making via consensus where at each iteration each robot needs to retain and transmit estimates of all robots’ actions [9, 10][11] requires sequential decision-making where all currently finalized actions need to be relayed to all robots in the network that have not finalized actions yet.
In more detail, the multi-robot coordination algorithm in [30], inspired by [9, 10, 11], achieves the best possible approximation bound for submodular maximization, namely, $1-1/e$. However, it may require tens of minutes to terminate in simulated tasks of $10$ robots even with no simulated communication delays [27]. The reason is that it requires a near-cubic number of iterations in the number of robots to converge (Table I). Similarly, for the Sequential Greedy (SG) algorithm [11], also known as Coordinate Descent, which is the gold standard in robotics and control for submodular task optimization [12, 2, 13, 15, 14, 1, 16, 17, 18, 19, 20], although it sacrifices some approximation performance to enable faster decision speed —achieving the bound $1/2$ instead of the bound $1-1/e$— has communication complexity that is cubic in the number of robots [28, Appendix II]. The reason is that it requires inter-robot messages that carry information about all the robots and a quadratic number of communication rounds over directed networks [31, Proposition 2].
@@ III Resource-Aware Distributed Submodular Optimization
$z_{i,t+1:t+T}\triangleq\{z_{i,t+1},\ldots,z_{i,t+T}\}$, $T$ is an action-planning look-ahead horizon, and $\mathbf{h}(\cdot|\cdot)$ is an information metric such as the conditional entropy [1] or the mutual information [15]. When $\mathbf{h}$ is conditional entropy and the process and sensor noises in eqs. 1, 3 and 2 are uncorrelated Gaussian, then eq. 4 is computable a priori given any action set $\{a_{i}\}_{i\,\in\,{\cal N}}$: $\mathbf{h}$ is then equal to the $\log\det$ volume of the Kalman filtering uncertainty over the horizon $T$ and is, thus, independent of the measurements’ realization [1].
Condition entropy, mutual information, and, more broadly, covering functions [1, 15, 44, 30, 45] are used to model active information-gathering tasks such as target tracking, SLAM, and surveillance [1, 15, 17]. These functions capture how much information is observed given the actions of all robots. They belong to a broader class of functions that satisfy the properties below (Definitions 1 and 2).
@@ V-A Centralization of Information
TABLE I: RAG vs. State-of-the-Art Distributed Submodular Maximization Algorithms. The state of the art is divided into algorithms that optimize (i) in the continuous domain, employing a continuous representation of $f$ [54], and (ii) in the discrete domain. For each algorithm, we present its decision time, split into computation and communication times, and suboptimality guarantee. We also specify the communication network topology that is required by each algorithm. The best performances for each metric are colored in blue. We assume for simplicity that $|{\cal V}_{i}|\,=|{\cal V}_{j}|,\forall i,j\in{\cal N}$.
@@ VI-B Comparison to the State of the Art
RAG requires at most $(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)$ communication time. In [37, 38], the agents need to communicate over ${\cal G}$ per the pre-specified information-access directed acyclic graph ${\cal G}_{\text{info}}$ per the rule in eq. 19 [37, Remark 3.2]. In Appendix IV, we identify a worst case where this rule results in $O(\tau_{c}\,|{\cal N}|^{2})$ communication time. This happens when each agent $i-1$ and agent $i$ in the decision sequence do not communicate directly, thus, the information of agent $i-1$ needs to be relayed to agent $i$ via other agents that form a connected communication path between the two. The method in [31], which introduces a depth-first search (DFS) procedure to determine the best agents’ ordering to run SG [11] over arbitrary (strongly) connected networks (instead of just line graphs), requires a worst-case communication time of $O(\tau_{c}\,|{\cal N}|^{3})$ for directed networks, and $O(\tau_{c}\,|{\cal N}|^{2})$ for undirected networks [28, Appendix II]. For some $\epsilon$, the method in [44] may require less communication time for running the algorithm per se than other methods, but an additional communication time of $O(\tau_{c}\,|{\cal V}_{i}|)$ is needed to distributively find a DAG that optimizes the algorithm’s approximation performance. The remaining algorithms require communication times proportional to $|{\cal N}|^{2}$ or more.
@@ VIII Summary and Concluding Remarks
We also plan the following algorithmic contributions. (i) RAG assumes synchronous communication. Although RAG can be modified to handle such cases, e.g., by instructing each robot to execute its action without first waiting for all other robots to select actions, its near-optimality guarantees provided in this paper may become invalid. Our future work will extend our theoretical and algorithmic analysis beyond the above limitations. (ii) We will also extend our results to handle effective task execution over long time horizons. For example, in collaborative mapping over long time horizons, the team needs to stay updated on the areas that have been mapped, such that the current plans do not repeat past actions. This is in addition to the focus of this paper that only the current plans among the robots do not overlap. To this end, we will handle network connectivity constraints. (iii) Finally, we will enhance our simulator by integrating the simulation of realistic communication channels and protocols towards communication-aware and -efficient coordination algorithms.
@@
$\leq(1+\kappa_{f})f({\cal A}^{{\sf RAG}})+\kappa_{f}\sum_{i\in{\cal N}}\left[f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}})\right]$ (27)
@@
where eq. 23 holds from telescoping the sums; eq. 24 holds from $f$ being submodular and $1-\kappa_{f}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}\setminus\{i\}}^{\sf RAG})}{f(a_{i}^{\sf RAG})}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}^{\sf OPT}\cup{\cal A}_{[i-1]}^{\sf RAG})}{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})}$ per the definition of $\kappa_{f}$; eq. 25 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily; eq. 26 holds from telescoping the sums; eq. 27 holds from $f$ being 2nd-order submodular; and eq. 28 holds from $f$ being submodular. Therefore, eq. 12 holds.
@@
where eq. 29 holds from eq. 8, and eq. 30 holds from [58, Lemma 2.1]. Combining eqs. 28 and 30, the lower bound in eq. 15 can be proved. ∎
@@
Definition 5 (Total curvature [9, 10]).

### v3

@@ I Introduction
Figure 1: Scenario of Collaborative Road Detection with 45 Aerial Robots. The drones are deployed in an unknown environment and tasked to collaboratively detect the roads and visually cover them. Particularly, the drones aim to collectively maximize the total covered road area in their top-view field of view (FOV). At each action coordination step, each drone chooses a few other drones to coordinate with, subject to its onboard communication and computation bandwidth constraints. The full collaborative autonomy pipeline is depicted in Fig. 2. (a)–(c) depict the task progress across time: (a) The drones start from their initial deployment points; (b) Progress after 3 replanning steps; (c) The drones have achieved maximal road coverage.
@@
The current approaches are either near-optimal but not scalable, or real-time but offer no performance guarantees. The near-optimal methods [9, 10] that are being used for active information gathering with multiple robots [11, 2, 12, 13, 1, 14, 15, 16, 17, 18, 19], scale cubically with the network size, resulting in impractical running times in mesh networks with tens of robots (up to tens of minutes per action coordination step in the presence of real-world communication delays). The state-of-the-art multi-robot motion planning methods are real-time [20, 21, 22, 23] but they are suitable in spatial exploration where robots are expected to operate far from each other. Then, the submodular structure can be ignored and opportunistic coordination is sufficient. We discuss the related work in extension in Section II.
@@
  • Time complexity: RAG’s decision-time scales linearly with the network size. Instead, the competitive near-optimal submodular optimization algorithms scale cubically with the network size. To this end, RAG introduces a resource-minimal distributed optimization paradigm that requires each robot to make decisions based only on locally observed information and information received from and about their neighbors only. In contrast, the current submodular optimization paradigms [9, 10], widely applied in robotics and control [11, 2, 12, 13, 1, 14, 15, 16, 17, 18, 19], that instead require the relay of information about all robots across the network.
@@ II Related Work
The current approximation paradigms for submodular maximization problem [9, 10] have been widely used in active information gathering with multiple robots [11, 2, 12, 13, 1, 14, 15, 16, 17, 18, 19] (see also the survey [5]). However, they may not scale. The reason is that their communication and computation complexities are superlinear in the number of robots, particularly, quadratic or cubic (Table I). Instead, we provide an algorithm with linear communication and computation complexity.
@@
The quadratic or higher complexity of [9, 10] over r2r networks is due to their coordination protocols: they instruct the robots to retain and relay information about all or most other robots in the network. Particularly, [9] requires iterative decision-making via consensus where at each iteration each robot needs to retain and transmit estimates of all robots’ actions. [10] requires sequential decision-making where all currently finalized actions need to be relayed to all robots in the network that have not finalized actions yet.
In more detail, the multi-robot coordination algorithm in [29], inspired by [9, 10], achieves the best possible approximation bound for submodular maximization, namely, $1-1/e$. However, it may require tens of minutes to terminate in simulated tasks of $10$ robots even with no simulated communication delays [26]. The reason is that it requires a near-cubic number of iterations in the number of robots to converge (Table I). Similarly, for the Sequential Greedy (SG) algorithm [10], also known as Coordinate Descent, which is the gold standard in robotics and control for submodular task optimization [11, 2, 12, 14, 13, 1, 15, 16, 17, 18, 19], although it sacrifices some approximation performance to enable faster decision speed —achieving the bound $1/2$ instead of the bound $1-1/e$— has communication complexity that is cubic in the number of robots [27, Appendix II]. The reason is that it requires inter-robot messages that carry information about all the robots and a quadratic number of communication rounds over directed networks [30, Proposition 2].
@@ III Resource-Aware Distributed Submodular Optimization
$z_{i,t+1:\tau}(a_{i})\triangleq\{z_{i,t+1}(a_{i}),\ldots,z_{i,\tau}(a_{i})\}$ denotes the measurements induced by $a_{i}$ up until $\tau$, $T$ is an action-planning look-ahead horizon, and $\mathbf{h}(\cdot|\cdot)$ is an information metric such as the conditional entropy [1] or the mutual information [14]. When $\mathbf{h}$ is conditional entropy and the process and sensor noises in eqs. 1, 3 and 2 are uncorrelated Gaussian, then eq. 4 is computable a priori given any action set $\{a_{i}\}_{i\,\in\,{\cal N}}$: $\mathbf{h}$ is then equal to the $\log\det$ volume of the Kalman filtering uncertainty over the horizon $T$ and is, thus, independent of the measurements’ realization [1].
Conditional entropy, mutual information, and, more broadly, covering functions [1, 14, 43, 29, 44] are used to model active information-gathering tasks such as target tracking, SLAM, and surveillance [1, 14, 16]. These functions capture how much information is observed given the actions of all robots. They belong to a broader class of functions that satisfy the properties below (Definitions 1 and 2).
@@ V-A Centralization of Information
TABLE I: RAG vs. State-of-the-Art Distributed Submodular Maximization Algorithms. The state of the art is divided into algorithms that optimize (i) in the continuous domain, employing a continuous representation of $f$ [53], and (ii) in the discrete domain. For each algorithm, we present its decision time, split into computation and communication times (see Section VI for definitions of $\tau_{f}$, $\tau_{c}$, and $\tau_{\#}$), and suboptimality guarantee. We also specify the communication network topology that is required by each algorithm. The best performances for each metric are in blue. We assume for simplicity that $|{\cal V}_{i}|\,=|{\cal V}_{j}|,\forall i,j\in{\cal N}$.
@@ VI-B Comparison to the State of the Art
RAG requires at most $(\tau_{c}+\tau_{\#})\,(|{\cal N}|-1)$ communication time. In [36, 37], the agents need to communicate over ${\cal G}$ per the pre-specified information-access directed acyclic graph ${\cal G}_{\text{info}}$ per the rule in eq. 19 [36, Remark 3.2]. In Appendix IV, we identify a worst case where this rule results in $O(\tau_{c}\,|{\cal N}|^{2})$ communication time. This happens when each agent $i-1$ and agent $i$ in the decision sequence do not communicate directly, thus, the information of agent $i-1$ needs to be relayed to agent $i$ via other agents that form a connected communication path between the two. The method in [30], which introduces a depth-first search (DFS) procedure to determine the best agents’ ordering to run SG [10] over arbitrary (strongly) connected networks (instead of just line graphs), requires a worst-case communication time of $O(\tau_{c}\,|{\cal N}|^{3})$ for directed networks, and $O(\tau_{c}\,|{\cal N}|^{2})$ for undirected networks [27, Appendix II]. For some $\epsilon$, the method in [43] may require less communication time for running the algorithm per se than other methods, but an additional communication time of $O(\tau_{c}\,|{\cal V}_{i}|)$ is needed to distributively find a DAG that optimizes the algorithm’s approximation performance. The remaining algorithms require communication times proportional to $|{\cal N}|^{2}$ or more.
@@ VIII Summary and Concluding Remarks
We also plan the following algorithmic contributions. (i) RAG assumes synchronous communication. Although RAG can be modified to handle such cases, e.g., by instructing each robot to execute its action without first waiting for all other robots to select actions, its near-optimality guarantees provided in this paper may no longer hold. Our future work will extend our theoretical and algorithmic analysis beyond the above limitations. (ii) We will also extend our results to handle effective task execution over long time horizons. For example, in collaborative mapping over long time horizons, the team needs to stay updated on the areas that have been mapped, such that the current plans do not repeat past actions. This is in addition to the focus of this paper that only the current plans among the robots do not overlap. To this end, we will handle network connectivity constraints. (iii) Finally, we will enhance our simulator by integrating the simulation of realistic communication channels and protocols towards communication-aware and -efficient coordination algorithms.
@@
$\leq(1+\kappa_{f})f({\cal A}^{{\sf RAG}})$
$\quad+\kappa_{f}\sum_{i\in{\cal N}}\left[f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}})\right]$ (27)
@@
where eq. 23 holds from telescoping the sums; eq. 24 holds from $f$ being submodular and
$1-\kappa_{f}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}\setminus\{i\}}^{\sf RAG})}{f(a_{i}^{\sf RAG})}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}^{\sf OPT}\cup{\cal A}_{[i-1]}^{\sf RAG})}{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})},$ (29)
per the definition of $\kappa_{f}$; eq. 25 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily; eq. 26 holds from telescoping the sums; eq. 27 holds from $f$ being 2nd-order submodular; and eq. 28 holds from $f$ being submodular. Therefore, eq. 12 holds.
@@
where eq. 30 holds from eq. 8, and eq. 31 holds from [57, Lemma 2.1]. Combining eqs. 28 and 31, the lower bound in eq. 15 can be proved. ∎
@@
Definition 5 (Total curvature [sviridenko2013optimal, 9]).